An angle measures 5 degrees more than 4 times its compliment. What is the measure of the angle. How is this problem solved?
Answer by nyc_function(2741)

Let A = the angle
Let (90 - A) = complement of the angle
A + 5 = 4(90 - A)
A + 5 = 360 - 4A
A + 4A = 360 - 5
5A = 355
A = 355/5
A = 71
The angle is 71 degrees.
Understand?